# Equivariant Ensembles and Regularization for Reinforcement Learning in Map-based Path Planning

### v2

@@ VI-A Setup
    Relative percentage deviation (RPD): Comparing the steps needed to solve an episode with a heuristic ([7]), showing the percentage difference. Lower is better.
@@ VI-C Performance
TABLE I: Quantitative comparison of the different agents tested on different sets of maps, showing the task-solved ratios and relative percentage deviation (RPD) to a heuristic for the in-distribution maps and the collection ratio (CR) for out-of-distribution maps. The Ens.+Reg.50M agent is the performance of Ens.+Reg. after 50M interaction steps.
This section compares the performance of each agent configuration on three different sets of maps: (1) the in-distribution maps that each agent encountered during training, (2) the in-distribution maps but rotated 90°, 180°, and 270°, and (3) 10 out-of-distribution maps shown in Figure 4(b) that no agent saw during training. Table I summarizes the results leading to the following observations. Each agent configuration consistently solves the in-distribution maps, with the Augment agent having the lowest RPD performance (highest RPD). As already seen in the training curves, the Ensemble and Regularized agents have similar performance compared with the baseline, and the combination has the best performance with a significant lead over the Baseline.

### v3

@@ VI-A Setup
    Relative deviation (RD): Comparing the steps needed to solve an episode with a heuristic ([7]), showing the difference. Lower is better.
@@ VI-C Performance
TABLE I: Quantitative comparison of the different agents tested on different sets of maps, showing the task-solved ratios and relative deviation (RD) to a heuristic for the in-distribution maps and the collection ratio (CR) for out-of-distribution maps. The Ens.+Reg.50M agent is the performance of Ens.+Reg. after 50M interaction steps. The numbers show the median value of the three agents per configuration, with the maximum deviation in superscript.
This section compares the performance of each agent configuration on three different sets of maps: (1) the in-distribution maps that each agent encountered during training, (2) the in-distribution maps but rotated 90°, 180°, and 270°, and (3) 10 out-of-distribution maps shown in Figure 4(b) that no agent saw during training. Table I summarizes the results leading to the following observations. Each agent configuration consistently solves the in-distribution maps, with the Augment agent having the lowest RD performance (highest RD). As already seen in the training curves, the Ensemble and Regularized agents have similar performance compared with the baseline, and the combination has the best performance with a significant lead over the Baseline.